B^{+} ∗ ∗ ∗ ∗ ∗

April 3, 2009

The B^{+} is the lightest charged meson containing a bottom quark. The mass and lifetime are well measured. It decays weakly and many of the decay modes are known, however these are not sufficient to saturated the total rate. We have therefore chosen to use the data from [?], together with more recent experimental results not included in [?], and some theoretical calculations to constructed the majority of the inclusive modes. Some additional modes are added to saturate some more exclusive measurements. The remainder of the decays are modelled using partonic decays.

The decays to D^{(∗)}D_{s}^{(∗)} are chosen to saturate the total rate 10.93% taken from [?] . Due to our choice of modes we have a lower value, 6.9%, for the D^{(∗)}D_{s} rate than the result from [?] . However given the error on this result and the total D^{(∗)}D_{s}^{(∗)} rate this is consistent with the experimental measurement.

The B^{+} has mass 5279 MeV and is unstable. The B^{+} has spin 0, charge 1 and is colour neutral. The B^{+} is a meson and is from the 1^{1}S_{0} multiplet.The properties of the particle and its antiparticle are taken to be independent. The decay length is 0.000501 m. The PDG code is 521.

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.2744378370  ∗∗  on

u, d̄, c̄, u Decay of B^{+} with branching ratio taken from EvtGen. This mode has minore modifications to ensure the total branching ratio is one and results from combining two modes from EvtGen due to the different modelling of partonic decays in Herwig and PYTHIA.

Herwig::WeakPartonic::WeakPartonic100

0.0740000000  ∗∗  on

c, s̄, c̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0617000000  ∗∗  on

D̄^{∗0}(2007), μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0617000000  ∗∗  on

D̄^{∗0}(2007), e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0330000000  ∗∗  on

(c̄d̄)_{0}, (uu)_{1} Decay of B^{+} with branching ratio taken from EvtGen. This mode is different as Herwig cannot hadronize diquarks containing a heavy quark. Therefore we have inserted an extra dd̄or uū pair with equal probability to produce a mode Herwig can handle. There are therefore two modes for the one baryon mode from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0278000000  ∗∗  on

D_{s}^{∗+}, D̄^{∗0}(2007)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0225000000  ∗∗  on

u, s̄, c̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0224000000  ∗∗  on

D̄^{0}, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0224000000  ∗∗  on

D̄^{0}, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0160000000  ∗∗  on

D̄^{∗0}(2007), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0159700000  ∗∗  on

D̄^{∗0}(2007), a_{1}^{+}(1260)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0150000000  ∗∗  on

D^{∗−}(2010), π^{0}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0134000000  ∗∗  on

ρ^{+}, D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0129000000  ∗∗  on

D̄^{0}, D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0124000000  ∗∗  on

D̄^{∗0}(2007), D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0111000000  ∗∗  on

D_{s}^{∗+}, D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0100000000  ∗∗  on

D̄^{∗0}(2007), D^{∗+}(2010), K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0098000000  ∗∗  on

D̄^{∗0}(2007), ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0098000000  ∗∗  on

D̄^{∗0}(2007), D_{s1}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0090000000  ∗∗  on

D̄′_{1}^{0}, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0090000000  ∗∗  on

(c̄s̄)_{0}, (uu)_{1} Decay of B^{+} with branching ratio taken from EvtGen. This mode is different as Herwig cannot hadronize diquarks containing a heavy quark. Therefore we have inserted an extra dd̄or uū pair with equal probability to produce a mode Herwig can handle. There are therefore two modes for the one baryon mode from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0090000000  ∗∗  on

D̄′_{1}^{0}, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0088700000  ∗∗  on

a_{1}^{+}(1260), D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0070000000  ∗∗  on

D̄^{0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::HQET

0.0070000000  ∗∗  on

D̄^{∗0}(2007), D^{∗0}(2007), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0056000000  ∗∗  on

D̄_{1}^{0}(2420), μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0056000000  ∗∗  on

D̄_{1}^{0}(2420), e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0050000000  ∗∗  on

D̄^{∗0}(2007), D^{∗+}(2010), K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0050000000  ∗∗  on

D̄^{∗0}(2007), D^{∗0}(2007), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0050000000  ∗∗  on

D̄^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0050000000  ∗∗  on

D̄^{0}, π^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0050000000  ∗∗  on

s̄, g, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0049000000  ∗∗  on

D̄_{0}^{∗0}, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0049000000  ∗∗  on

D̄_{0}^{∗0}, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0049000000  ∗∗  on

D̄^{0}, D^{∗+}(2010), K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0047000000  ∗∗  on

D̄^{0}, D^{∗0}(2007), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0046000000  ∗∗  on

D̄^{∗0}(2007), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0042000000  ∗∗  on

D̄_{2}^{∗0}(2460), D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0041000000  ∗∗  on

D_{s1}^{+}, D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0040000000  ∗∗  on

D̄_{2}^{∗0}(2460), D_{s}^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0040000000  ∗∗  on

c, d̄, c̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0040000000  ∗∗  on

c, s̄, ū, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0038000000  ∗∗  on

D̄_{2}^{∗0}(2460), ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0037000000  ∗∗  on

D_{s}^{∗+}, D^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0036000000  ∗∗  on

D_{s}^{+}, D^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0033000000  ∗∗  on

D_{s}^{+}, D^{−}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0033000000  ∗∗  on

D_{s}^{+}, D̄^{0}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0033000000  ∗∗  on

D_{s}^{∗+}, D^{−}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0033000000  ∗∗  on

D_{s}^{∗+}, D̄^{0}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0031000000  ∗∗  on

D̄^{∗0}(2007), D^{+}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0030000000  ∗∗  on

D̄_{2}^{∗0}(2460), μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0030000000  ∗∗  on

D̄_{2}^{∗0}(2460), e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0030000000  ∗∗  on

u, d̄, ū, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0025000000  ∗∗  on

D̄^{0}, D^{+}, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0025000000  ∗∗  on

D̄^{∗0}(2007), D^{+}, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0025000000  ∗∗  on

D̄^{0}, D^{∗+}(2010), K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0025000000  ∗∗  on

D̄^{0}, D^{0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0025000000  ∗∗  on

D̄^{∗0}(2007), D^{0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0025000000  ∗∗  on

D̄^{0}, D^{∗0}(2007), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0024000000  ∗∗  on

D̄_{1}^{0}(2420), D_{s}^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0022000000  ∗∗  on

D̄′_{1}^{0}, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0022000000  ∗∗  on

u, ū, s̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0021000000  ∗∗  on

D̄^{0}, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0020500000  ∗∗  on

u, ū, d̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0020000000  ∗∗  on

D̄_{2}^{∗0}(2460), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0020000000  ∗∗  on

D̄′_{1}^{0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0020000000  ∗∗  on

D^{−}, ρ^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0020000000  ∗∗  on

D^{−}, π^{0}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0020000000  ∗∗  on

D^{∗−}(2010), ρ^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0019480000  ∗∗  on

μ^{+}, ν_{μ}, ū, u Decay of B^{+} with branching ratio taken from EvtGen, this partonic mode replaces the specialised model in EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0019480000  ∗  on

e^{+}, ν_{e}, ū, u Decay of B^{+} with branching ratio taken from EvtGen, this partonic mode replaces the specialised model in EvtGen.

Herwig::WeakPartonic::WeakPartonic100

0.0018000000  ∗∗  on

J/ψ(1S), K_{1}^{+}(1270)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0018000000  ∗∗  on

D_{s}^{+}, D̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0018000000  ∗∗  on

D_{s}^{∗+}, D̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0018000000  ∗∗  on

D̄^{∗0}(2007), D^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0018000000  ∗∗  on

d, d̄, s̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0017000000  ∗∗  on

D̄^{0}, D^{+}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0016000000  ∗∗  on

D̄^{0}, D_{s0}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0016000000  ∗∗  on

D̄^{∗0}(2007), D_{s0}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0015000000  ∗  on

D̄^{0}, D^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0015000000  ∗∗  on

D^{+}, D^{∗−}(2010), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0015000000  ∗∗  on

D^{∗+}(2010), D^{∗−}(2010), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0015000000  ∗∗  on

s, s̄, s̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0014300000  ∗∗  on

J/ψ(1S), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0013000000  ∗∗  on

D̄_{0}^{∗0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0013000000  ∗∗  on

D̄_{1}^{0}(2420), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0012500000  ∗∗  on

D^{∗−}(2010), π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0012000000  ∗  on

K^{∗+}, η_{c} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0012000000  ∗∗  on

D̄′_{1}^{0}, D_{s}^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0012000000  ∗∗  on

D̄_{1}^{0}(2420), D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0010600000  ∗∗  on

J/ψ(1S), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0010400000  ∗∗  on

D̄^{∗0}(2007), D^{∗+}(2010)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0010200000  ∗∗  on

D^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0010000000  ∗∗  on

D^{∗+}(2010), D^{∗−}(2010), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0010000000  ∗∗  on

D̄_{1}^{0}(2420), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0009100000  ∗∗  on

D̄_{0}^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0009000000  ∗  on

η_{c}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0009000000  ∗∗  on

s, s̄, d̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0008000000  ∗∗  on

D_{s}^{+}, D̄^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0008000000  ∗∗  on

D_{s}^{∗+}, D̄^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0008000000  ∗∗  on

D̄_{2}^{∗0}(2460), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0007700000  ∗∗  on

D̄^{∗0}(2007), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0007500000  ∗∗  on

D̄′_{1}^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0007000000  ∗∗  on

D̄_{1}^{0}(2420), ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0007000000  ∗∗  on

d, d̄, d̄, u Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::WeakPartonic::WeakPartonic0

0.0006300000  ∗∗  on

ψ(2S), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0006100000  ∗∗  on

K^{∗+}, D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0006000000  ∗∗  on

ψ(2S), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0006000000  ∗∗  on

D̄′_{1}^{0}, D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0005700000  ∗  on

χ_{c1}(1P), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

J/ψ(1S), K_{2}^{∗+}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0005000000  ∗  on

ψ(3370), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D^{+}, D^{−}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D^{∗+}(2010), D^{−}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D^{+}, D^{−}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D^{∗+}(2010), D^{−}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D^{+}, D^{∗−}(2010), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D̄^{0}, π^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D̄^{∗0}(2007), π^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D̄^{∗0}(2007), ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0005000000  ∗∗  on

D̄^{∗0}(2007), π^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0005000000  ∗∗  on

D̄^{∗0}(2007), ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0005000000  ∗∗  on

D̄^{∗0}(2007), π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0004800000  ∗  on

K^{∗+}, η_{c}(2S)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0004800000  ∗  on

ψ(3370), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0004100000  ∗∗  on

D̄^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0004000000  ∗∗  on

ψ(2S), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0004000000  ∗∗  on

ψ(2S), K_{1}^{+}(1270)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0004000000  ∗  on

K^{∗+}, χ_{c0}(1P)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0004000000  ∗  on

χ_{c1}(1P), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0004000000  ∗  on

χ_{c1}(1P), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0003600000  ∗  on

η_{c}(2S), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0003600000  ∗  on

D̄^{∗0}(2007), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0003350000  ∗∗  on

D^{∗+}(2010), D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0003350000  ∗∗  on

D̄^{∗0}(2007), D^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0003118000  ∗∗  on

s̄, u, γ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::BtoSGamma::BtosgammaKagan

0.0003000000  ∗  on

χ_{c0}(1P), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0003000000  ∗  on

ψ(3370), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0003000000  ∗  on

ψ(3370), K_{1}^{+}(1270)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002900000  ∗  on

χ_{c1}(1P), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002700000  ∗∗  on

D^{+}, D̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0002000000  ∗∗  on

J/ψ(1S), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

ψ(2S), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

ψ(2S), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0002000000  ∗  on

η_{c}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

η_{c}, K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0002000000  ∗  on

χ_{c0}(1P), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

χ_{c0}(1P), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0002000000  ∗  on

χ_{c1}(1P), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

χ_{c1}(1P), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0002000000  ∗  on

χ_{c1}(1P), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0002000000  ∗  on

χ_{c2}(1P), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

χ_{c2}(1P), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0002000000  ∗  on

ψ(3370), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0002000000  ∗  on

ψ(3370), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001450000  ∗∗  on

ω, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0001450000  ∗∗  on

ρ^{0}, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0001450000  ∗∗  on

ω, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0001450000  ∗  on

ρ^{0}, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0001000000  ∗∗  on

J/ψ(1S), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0001000000  ∗∗  on

J/ψ(1S), K_{1}^{+}(1400)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0001000000  ∗  on

ψ(2S), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

ψ(2S), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

η_{c}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0001000000  ∗  on

η_{c}, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

η_{c}, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

χ_{c0}(1P), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0001000000  ∗  on

χ_{c0}(1P), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

χ_{c0}(1P), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

χ_{c2}(1P), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0001000000  ∗  on

χ_{c2}(1P), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

χ_{c2}(1P), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

ψ(3370), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0001000000  ∗  on

ψ(3370), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000930000  ∗  on

τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000900000  ∗  on

J/ψ(1S), ϕ, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000840000  ∗∗  on

η′, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000840000  ∗∗  on

η, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000840000  ∗∗  on

η′, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000840000  ∗∗  on

η, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000800000  ∗  on

η_{c}(2S), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000800000  ∗  on

η_{c}(2S), K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000780000  ∗  on

η′, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000720000  ∗∗  on

π^{0}, μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000720000  ∗∗  on

π^{0}, e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000700000  ∗  on

K_{0}^{∗0}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000600000  ∗∗  on

J/ψ(1S), ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0000500000  ∗  on

a_{1}^{+}(1260), a_{1}^{0}(1260)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000500000  ∗  on

b_{1}^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000500000  ∗  on

b_{1}^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000500000  ∗  on

η_{c}(2S), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000500000  ∗  on

η_{c}(2S), K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000500000  ∗  on

η_{c}(2S), K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000480000  ∗  on

K^{+}, K^{0}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000460000  ∗∗  on

a_{1}^{0}(1260), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000420000  ∗∗  on

ω, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0000420000  ∗∗  on

ρ^{0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZVector

0.0000403000  ∗  on

K^{∗+}, γ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000400000  ∗  on

a_{1}^{+}(1260), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000400000  ∗  on

a_{1}^{0}(1260), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000400000  ∗∗  on

J/ψ(1S), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::ScalarMesonFactorized::ISGW2Hadronic

0.0000300000  ∗∗  on

π^{0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000280000  ∗  on

D_{s}^{∗+}, ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000280000  ∗  on

ρ^{0}, D_{s}^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000270000  ∗∗  on

b_{1}^{0}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000260000  ∗  on

K^{+}, K^{−}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000260000  ∗  on

π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000240000  ∗  on

K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000240000  ∗  on

K^{∗+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000230000  ∗∗  on

f_{1}(1420), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000230000  ∗∗  on

f_{1}(1285), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000200000  ∗  on

a_{1}^{+}(1260), K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗  on

a_{1}^{0}(1260), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗  on

D_{s}^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗  on

D_{s}^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗∗  on

f′_{2}(1525), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗∗  on

f_{2}(1270), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000200000  ∗  on

χ_{c2}(1P), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000200000  ∗  on

χ_{c2}(1P), K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000180000  ∗∗  on

η′, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000160000  ∗  on

π^{+}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000150000  ∗  on

ρ^{+}, ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000130000  ∗  on

ω, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000130000  ∗∗  on

h′_{1}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000130000  ∗∗  on

h_{1}, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000120000  ∗  on

ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000120000  ∗  on

K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000120000  ∗∗  on

η, τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::BZScalar

0.0000110000  ∗  on

ϕ, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000110000  ∗  on

D^{+}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K_{1}^{0}(1400), K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{0}, K̄_{1}(1400)^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000100000  ∗  on

K^{+}, K_{0}^{∗0}(1430), K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K̄_{0}^{∗0}(1430), K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{0}, K^{0}, K̄_{0}^{∗0}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{+}, K̄_{0}^{∗0}(1430), π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K_{0}^{∗0}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{+}, K_{0}^{∗−}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K_{0}^{∗+}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ϕ, K_{0}^{∗+}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K_{0}^{∗0}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{0}, K̄^{0}, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{0}, K̄^{0}, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{−}, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{−}, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K^{0}, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K^{+}, f_{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K^{+}, ω Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K^{+}, ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{0}, K^{∗+}, K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{0}, K̄^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{∗0}, K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{∗−}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{−}, K^{+}, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ϕ, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ϕ, K^{+}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ϕ, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ϕ, K^{0}, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K^{+}, f_{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K^{+}, ω Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000100000  ∗  on

ϕ, K^{+}, ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), K^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{0}(1260), K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{0}(1260), K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{∗+}, f_{0}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{∗0}, f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{0}, π^{0}, K_{0}^{∗+}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{0}, K_{0}^{∗0}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{+}, K_{0}^{∗0}(1430), π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{−}, K_{0}^{∗+}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K_{0}^{∗+}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K_{0}^{∗+}(1430), π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K_{0}^{∗0}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K_{0}^{∗0}(1430), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ω, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{0}, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{0}, K^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{−}, K^{0}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{−}, K^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{+}, K^{+}, π^{−}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{+}, K^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{+}, K^{0}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{0}, π^{0}, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{−}, π^{+}, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{0}, K^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000100000  ∗  on

π^{+}, π^{−}, K^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ω, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, K^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{−}, K^{∗+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K^{∗+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, a_{0}^{+}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{−}, a_{0}^{+}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, a_{0}^{0}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, f_{0}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{−}(1260), π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{0}(1260), π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), f_{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), ω Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, ρ^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, ρ^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K_{0}^{∗0}(1430), K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K̄_{0}^{∗0}(1430), K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000100000  ∗  on

K^{+}, K^{+}, K_{0}^{∗−}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{−}, K^{+}, K_{0}^{∗+}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K_{0}^{∗+}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{0}, K̄^{0}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{0}, K̄^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{−}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{+}, K^{−}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

K^{0}, K^{∗+}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K^{0}, K̄^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K^{∗0}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{+}, K^{∗−}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{−}, K^{+}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ϕ, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

b_{1}^{+}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

b_{1}^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{+}(1260), K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

a_{1}^{0}(1260), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

K^{∗+}, f_{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{0}, π^{0}, K_{0}^{∗+}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{+}, π^{0}, K_{0}^{∗0}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K_{0}^{∗+}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K_{0}^{∗0}(1430)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{0}, π^{0}, π^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{−}, π^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{0}, π^{0}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{+}, π^{+}, π^{−}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{−}, K^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000100000  ∗  on

ρ^{+}, K^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{0}, π^{0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{−}, π^{+}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{+}, π^{0}, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

f_{0}(980), π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, a_{0}^{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, f_{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

π^{+}, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

π^{−}, π^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000100000  ∗  on

ρ^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{+}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ω, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000100000  ∗  on

ρ^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000091000  ∗  on

ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000090000  ∗  on

ϕ, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000086000  ∗  on

ρ^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000085000  ∗  on

f_{0}(980), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000060000  ∗  on

D^{∗+}(2010), K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000059000  ∗  on

ω, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000055000  ∗  on

π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000052000  ∗  on

ρ^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000051000  ∗  on

ω, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{+}, K_{1}^{0}(1400), K̄^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{+}, K^{0}, K̄_{1}(1400)^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{+}, K_{0}^{∗0}(1430), K̄^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000050000  ∗  on

K^{+}, K̄_{0}^{∗0}(1430), K^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{+}, K^{+}, K_{0}^{∗−}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{−}, K^{+}, K_{0}^{∗+}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ϕ, K_{0}^{∗+}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{0}, K̄^{0}, K^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{+}, K^{−}, K^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ϕ, K^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{∗0}, a_{0}^{+}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{∗+}, a_{0}^{0}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{∗+}, f_{0}(980), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

π^{0}, π^{0}, K_{0}^{∗+}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{+}, π^{0}, K_{0}^{∗0}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ρ^{0}, K_{0}^{∗+}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, K_{0}^{∗0}(1430), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ω, K^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ρ^{0}, K^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ρ^{+}, K^{0}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{0}, η, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{−}, π^{+}, K^{∗+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{+}, η, K^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ω, K^{∗+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{0}, K^{∗+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, K^{∗0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{0}, a_{0}^{+}(980), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, a_{0}^{0}(980), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, f_{0}(980), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

a_{1}^{+}(1260), π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

a_{1}^{0}(1260), π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000050000  ∗  on

ω, ρ^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, ρ^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{0}, K̄^{0}, K^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

K^{+}, K^{−}, K^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ϕ, K^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{∗0}, a_{0}^{+}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

K^{∗+}, a_{0}^{0}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

π^{0}, π^{0}, η, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{+}, π^{−}, η, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{+}, η, π^{0}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ω, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ω, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{0}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{+}, K^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

π^{0}, η, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

π^{+}, η, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

π^{+}, η, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

π^{−}, π^{+}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000050000  ∗  on

ρ^{+}, π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ω, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000050000  ∗  on

ρ^{0}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000049000  ∗  on

K^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000048000  ∗  on

η, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000042000  ∗  on

η′, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000040000  ∗∗  on

a_{0}^{0}(980), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000030000  ∗  on

ϕ, ϕ, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000030000  ∗  on

ϕ, ϕ, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000030000  ∗  on

K^{∗+}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000026000  ∗  on

η, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000023000  ∗  on

f_{2}(1270), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000022000  ∗  on

ρ^{0}(1450), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{+}, K_{1}^{0}(1400), K̄^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{0}, K̄_{1}(1400)^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K̄^{0}, K̄^{0}, K_{1}^{+}(1400), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{0}, K^{0}, K̄_{1}(1400)^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{0}, K̄^{0}, K_{1}^{+}(1400), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{0}, K̄^{0}, K_{1}^{0}(1400), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{+}, K̄_{1}(1400)^{0}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{+}, K_{1}^{−}(1400), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{−}, K^{+}, K_{1}^{+}(1400), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{−}, K^{+}, K_{1}^{0}(1400), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ϕ, K_{1}^{+}(1400), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ϕ, K_{1}^{0}(1400), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{+}, K_{0}^{∗0}(1430), K̄^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K̄_{0}^{∗0}(1430), K^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{+}, K_{0}^{∗−}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{−}, K^{+}, K_{0}^{∗+}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ϕ, K_{0}^{∗+}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{0}, K̄^{0}, K^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{−}, K^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ϕ, K^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{∗+}, f_{0}(980), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

π^{0}, π^{0}, K_{0}^{∗+}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, π^{0}, K_{0}^{∗0}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{0}, K_{0}^{∗+}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000020000  ∗  on

ρ^{+}, K_{0}^{∗0}(1430), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

π^{0}, π^{0}, π^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, π^{−}, π^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, π^{−}, π^{+}, K^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, π^{0}, π^{0}, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, π^{+}, π^{−}, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ω, K^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{0}, K^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{+}, K^{0}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{0}, η′, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

η, η, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{−}, π^{+}, K^{∗+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, η′, K^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ω, K^{∗+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{0}, K^{∗+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, K^{∗0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

f_{0}(980), π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

f_{0}(980), π^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{0}, a_{0}^{+}(980), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, a_{0}^{0}(980), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, f_{0}(980), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

a_{1}^{+}(1260), π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

a_{1}^{0}(1260), π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{+}, π^{+}, π^{−}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{−}, π^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ω, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{0}, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000020000  ∗  on

ρ^{0}, π^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ω, ρ^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, ρ^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{−}, K^{+}, K_{1}^{+}(1400)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ϕ, K_{1}^{+}(1400)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{0}, K̄^{0}, K^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

K^{+}, K^{−}, K^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ϕ, K^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

π^{0}, π^{0}, η′, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{0}, η, η, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, η, η, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ω, K^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{0}, K^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, K^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

π^{0}, η′, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

η, η, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{0}, a_{0}^{+}(980)

Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

π^{+}, η′, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{+}, η, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

π^{−}, π^{+}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000020000  ∗  on

ρ^{+}, π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{+}, η, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ω, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

ρ^{0}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K_{0}^{∗+}(1430), K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K_{0}^{∗+}(1430), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000020000  ∗  on

K̄_{0}^{∗0}(1430), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000020000  ∗∗  on

f_{0}(1370), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::SemiLeptonicScalar::ISGW2

0.0000020000  ∗∗  on

f_{0}(980), τ^{+}, ν_{τ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000018350  ∗  on

K^{∗+}, e^{+}, e^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000013800  ∗  on

K^{∗+}, μ^{+}, μ^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, K̄^{0}, K^{+}, K^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{0}, K̄^{0}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{0}, K̄^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{+}, K^{−}, K^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{+}, K^{−}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K̄^{0}, K^{+}, K^{∗0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{0}, K̄^{0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{+}, K^{−}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, K̄^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, K^{+}, K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, K^{0}, K̄^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, K^{+}, K^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K̄^{0}, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{0}, K̄^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{0}, K̄^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K^{+}, K^{−}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, ϕ, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K̄^{0}, K̄^{0}, K_{1}^{+}(1400), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K̄^{0}, K_{1}^{+}(1400), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, K_{1}^{+}(1400), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, K_{1}^{+}(1400), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{−}, K^{+}, K_{1}^{+}(1400), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{−}, K^{+}, K_{1}^{+}(1400), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K_{1}^{+}(1400), η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000010000  ∗  on

ϕ, K_{1}^{+}(1400), η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗+}, K̄^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K̄^{∗0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{∗0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{∗0}, K̄^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{∗−}, K^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{∗+}, K^{−}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{∗−}, K^{0}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{∗0}, K^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, f_{0}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, ω, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, ρ^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, ρ^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, ρ^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, ω, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, ω, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K^{−}, ρ^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, π^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, ρ^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, ω, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, ρ^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000010000  ∗  on

K̄^{∗0}, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗0}, K̄^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{0}, η, η′, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{0}, π^{0}, η′, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{0}, η, η, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{0}, π^{0}, η, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, π^{−}, η′, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, π^{−}, η, K^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, η, η, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, η, π^{0}, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, π^{+}, π^{−}, K^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, π^{+}, π^{−}, K^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

η, η′, K^{∗+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

a_{0}^{−}(980), π^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

a_{0}^{+}(980), π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

a_{0}^{0}(980), π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

f_{0}(980), π^{+}, π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

f_{0}(980), π^{+}, π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, η′, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, η, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, π^{0}, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{−}, π^{+}, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{−}, π^{+}, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{−}, π^{+}, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{+}, π^{0}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{+}, η, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{+}, π^{0}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{+}, π^{+}, π^{−}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000010000  ∗  on

ρ^{+}, π^{+}, π^{−}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{−}, π^{+}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{−}, π^{+}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ω, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ω, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{0}, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ρ^{0}, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, K̄^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, K^{+}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ϕ, K̄^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗+}, K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K̄^{∗0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{0}, K̄^{∗0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗0}, K̄^{0}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗−}, K^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗+}, K^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K̄^{0}, K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, π^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{0}, K̄^{0}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K^{+}, K^{−}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

K̄^{0}, K^{+}, ω Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K̄^{0}, K^{+}, ρ^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{0}, K̄^{0}, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K^{−}, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, ρ^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K̄^{∗0}, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000010000  ∗  on

π^{0}, η, η′, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

ω, K^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ρ^{0}, K^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

η, η′, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{−}(980), π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{+}(980), π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{0}(980), π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, η′, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000010000  ∗  on

π^{+}, K^{0}, K̄^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, η′, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, η, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, η, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, η, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, η, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, K^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, K^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

π^{+}, K^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K̄^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K̄^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K̄^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{+}, K^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{+}(980), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000010000  ∗  on

a_{0}^{0}(980), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

a_{0}^{+}(980), K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

f_{0}(980), π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ρ^{0}(1450), K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ρ^{+}, K^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K̄^{0}, K^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ϕ, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ω, K^{∗+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

ρ^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000010000  ∗  on

K^{∗+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005740  ∗  on

K^{+}, μ^{+}, μ^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005740  ∗  on

K^{+}, e^{+}, e^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005400  ∗  on

D^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005400  ∗  on

D^{∗+}(2010), π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005000  ∗  on

K^{∗+}, K̄^{0}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K^{+}, K̄^{∗0}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K̄^{0}, K^{+}, ρ^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K^{0}, K̄^{0}, ρ^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K^{+}, K^{−}, ρ^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

ϕ, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

ϕ, ρ^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005000  ∗  on

K̄^{∗0}, K^{∗+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005000  ∗  on

K^{∗+}, K̄^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005000  ∗  on

K^{+}, K̄^{∗0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000005000  ∗  on

K̄^{0}, K^{+}, π^{0}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K^{0}, K̄^{0}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

K^{+}, K^{−}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000005000  ∗  on

ϕ, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000002000  ∗  on

K^{∗+}, K̄^{0}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{+}, K̄^{∗0}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K̄^{0}, K^{+}, π^{0}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K̄^{0}, K^{+}, π^{+}, π^{−}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{0}, K̄^{0}, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{0}, K̄^{0}, π^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{+}, K^{−}, π^{+}, π^{0}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{+}, K^{−}, π^{+}, π^{−}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{0}, K^{−}, π^{+}, π^{+}, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K̄^{0}, K^{+}, ρ^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{0}, K̄^{0}, ρ^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{+}, K^{−}, ρ^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

ϕ, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

ϕ, ρ^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000002000  ∗  on

K̄^{∗0}, K^{∗+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000002000  ∗  on

K^{∗+}, K̄^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000002000  ∗  on

K^{+}, K̄^{∗0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000002000  ∗  on

K̄^{0}, K^{+}, π^{0}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{0}, K̄^{0}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

K^{+}, K^{−}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000002000  ∗  on

ϕ, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000001900  ∗  on

K^{∗+}, τ^{+}, τ^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000001300  ∗  on

K^{+}, τ^{+}, τ^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000001000  ∗  on

K̄^{0}, K^{+}, π^{0}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K̄^{0}, K^{+}, π^{0}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K̄^{0}, K^{+}, π^{+}, π^{−}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K̄^{0}, K^{+}, π^{+}, π^{−}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{0}, K̄^{0}, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

Table 1: The decay modes of the B^{+} (continues).

Branching  Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles
0.0000001000  ∗  on

K^{0}, K̄^{0}, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{+}, K^{−}, π^{+}, η′, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{+}, K^{−}, π^{+}, η, π^{0} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{0}, K^{−}, π^{+}, π^{+}, η′ Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{0}, K^{−}, π^{+}, π^{+}, η Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Mambo::Mambo

0.0000001000  ∗  on

K^{−}, π^{+}, π^{+} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000001000  ∗  on

K^{+}, K^{+}, π^{−} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000000000  ∗  on

e^{+}, ν_{e} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

0.0000000000  ∗  on

μ^{+}, ν_{μ} Decay of B^{+} with branching ratio taken from EvtGen.

Herwig::Hw64::DecayME0

Table 1: The decay modes of the B^{+}.

The B^{+} decay modes are given in Table 1 and the total branching ratio is 0.99999999999999. There is no mass generator for the B^{+}. There is no width generator for the B^{+}. The particle was checked by Peter Richardson at 8:46:58 on the 16th of May 2008. The most recently changed decay mode was

modified at 8:44:59 on the 16th of May 2008. The particle data was last modified at 14:33:4 on the 15th of May 2008.